Write a balanced net ionic equation for MgCl2(aq) + Li2SO4(aq) ---> MgSO4(s) + 2LiCl(aq)
Answer #1

MgCl2 (aq) + Li2SO4 (aq) = MgSO4 (s) + 2LICl (aq)

Complete ionic equation is -

Mg2+ (aq) + 2Cl- (aq) + 2Li+ (aq) + SO42-=  MgSO4 (s) + 2LI+ + 2Cl- (aq)

So, Net Ionic equation is -

Mg2+ (aq) + SO42- (aq) =  MgSO4 (s)